OpenStudy (anonymous):

Separate 60 into two parts so that 3 times the smaller added to 6 more than 6 times the smaller = 60. help?

hartnn (hartnn):

let one part be x then other part = 60-x ok? then verify next part of your question....is it smaller at both place?

OpenStudy (anonymous):

I still don't understand..

OpenStudy (lgbasallote):

3 times the smaller would be 3x 6 more than 6 times the smaller is 6x + 6 it says the sum of that is 60 so 3x + 6x + 6 = 60 combine like terms 9x + 6 = 60 subtract 6 from both sides 9x = 54 do you know what to do next now?

OpenStudy (anonymous):

Divide 54 by 9 which would be 6.. Am I right ?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Is that it ?

OpenStudy (lgbasallote):

it is

OpenStudy (anonymous):

oh ok.. thank yous sooo much ^.^

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

Wow fate you do ogburn ??